OpenStudy (anonymous):

There is a question as following \[0,\overline{1101}=\sum_{k=0}^{\infty}\frac{1}{2^{4k+1}}+\frac{1}{2^{4k+2}}+\frac{1}{2^{4k+4}}\] but i dont understand this following part where the nominators coming \[=\sum_{k=0}^{\infty}\frac{2^{2}.2^{4}+2.2^{4}+2.2^{2}}{2^{4k}.2^{7}}\]

OpenStudy (amistre64):

\[\frac{this}{that}\] well, latex is working

OpenStudy (amistre64):

what does the 0,overline mean?

OpenStudy (anonymous):

sorry my internet connection is bad, i didnt got chance to make it better, now its a little bit better, i will correct it

OpenStudy (amistre64):

\[0,\overline{1101} =\sum_{k=0}^{\infty}\frac{1}{2^{4k+1}} +\frac{1} {(2^{4k+1})^{2} } +\frac{1}{(2^{4k+1})^{4}}\]

OpenStudy (amistre64):

the tex doesnt do a double ^ very well 2^x^k fails \[2^x^k\]

OpenStudy (anonymous):

very near but its not that just a moment amistre

ganeshie8 (ganeshie8):

*